### v1

# Emergent Symmetry and Phase Transitions on the Domain Wall of $\mathbb{Z}_{2}$ Topological Orders

Abstract
The one-dimensional (1D) domain wall of 2D $\mathbb{Z}_{2}$ topological orders is studied theoretically. The Ising domain wall model is shown to have an emergent SU(2)1 conformal symmetry because of a hidden nonsymmorphic octahedral symmetry. While a weak magnetic field is an irrelevant perturbation to the bulk topological orders, it induces a domain wall transition from the Tomonaga-Luttinger liquid to a ferromagnetic order, which spontaneously breaks the anomalous $\mathbb{Z}_{2}$ symmetry and the time-reversal symmetry on the domain wall. Moreover, the gapless domain wall state also realizes a 1D topological quantum critical point between a $\mathbb{Z}_{2}^{T}$-symmetry-protected topological phase and a trivial phase, thus demonstrating the holographic construction of topological transitions.
Introduction.— Topological phases often have nontrivial gapless boundary states, which bear valuable information of the bulk topological states via the bulk-boundary correspondence. The chiral edge states of quantum Hall insulators (QHIs) [1, 2, 3] and the helical edge states of quantum spin Hall insulators (QSHIs) [4, 5] are prominent examples. The chiral edge states of QHIs are robust against generic local perturbations as long as the bulk energy gap is not closed, and are characterized by a nonzero thermal Hall conductance [6, 7]. In contrast, the helical edge states of QSHIs are protected only when the time-reversal symmetry [8, 9, 10] or the spin rotation symmetry [11] is preserved, which prohibits the backscattering of electrons at the edge.
Nonchiral topological phases with vanishing thermal Hall conductance may have gapped boundary states, where a group of bosonic topological excitations condense at the boundary and render all other topological excitations confined because of their nontrivial mutual statistics with the condensed ones [12, 13, 14, 15, 16, 17]. Gapped domain walls between topological orders can also be constructed if topological excitations can tunnel through the domain wall in a way compatible with their braiding statistics in the bulk topological orders [18].
Gapless domain walls of nonchiral topological orders are more elusive. Recently, a systematic construction of one-dimensional (1D) gapless domain walls of 2D topological orders was proposed [19]. In particular, for the domain wall between $\mathbb{Z}_{2}$ topological orders, i.e., the toric code [20, 21] and the double semion models [21], the low-energy spectrum was calculated numerically and found to be consistent with the SU(2)1 Wess-Zumino-Witten (WZW) conformal field theory [19], even though the domain wall model does not have any continuous symmetry. An argument based on the anyon condensation scenario was proposed for the emergent SU(2)1 conformal symmetry [19], but a microscopic understanding is still lacking. In addition, it is also desirable to find out the general domain wall phase diagram and particularly the possible transitions to gapped domain wall states.
In this work, we study the emergent symmetry and the phase diagram on the domain wall of $\mathbb{Z}_{2}$ topological orders. We find an exact unitary transformation that maps the domain wall model into a deformed Heisenberg chain and thus identify a hidden nonsymmorphic octahedral symmetry. The low-energy effective theory is given by the SU(2)1 WZW model, because all relevant perturbations are prohibited by the nonsymmorphic octahedral symmetry. We then consider a more generic domain wall perturbed by a nonzero magnetic field of the bulk topological orders, which induces an Ising interaction on the domain wall model. It drives the domain wall transitions from the gapless Tomonaga-Luttinger liquid (TLL) into gapped phases with ferromagnetic (FM) or antiferromagnetic (AF) orders, which spontaneously break the time-reversal symmetry and the anomalous $\mathbb{Z}_{2}$ symmetry inherited from the bulk $\mathbb{Z}_{2}$ gauge structure. Moreover, the gapless domain wall state corresponds to a 1D quantum critical point (QCP) between a $\mathbb{Z}_{2}^{T}$-symmetry-protected topological (SPT) phase and a trivial phase, hence it is an example of the holographic construction of topological QCPs as gapless domain walls of one-higher-dimensional topological phases [22, 23].
Emergent SU(2)1 conformal symmetry.— The 1D domain wall between the toric code and the double semion topological orders is described by the Ising domain wall model (IDWM) [19],
$H_{0}=-\sum_{l}\sigma^{z}_{l-1}\sigma^{x}_{l}\sigma^{z}_{l+1}-\sum_{l}\sigma^{y}_{l},$ (1)
where $\sigma^{x}_{l}$, $\sigma^{y}_{l}$ and $\sigma^{z}_{l}$ are the Pauli operators on site $l$. The IDWM has an anomalous $\mathbb{Z}_{2}$ self-dual symmetry $\mathcal{S}$, which is inherited from the bulk $\mathbb{Z}_{2}$ gauge structure [24, 19],
$\mathcal{S}=\prod_{l}\sigma^{x}_{l}\prod_{l}\exp\Big{(}\frac{i\pi}{4}(\sigma^{z}_{l}\sigma^{z}_{l+1}-\sigma^{z}_{l}-1)\Big{)},$ (2)
which exchanges the two terms in Eq. (1). $H_{0}$ also has an antiunitary $\mathbb{Z}_{2}^{T}$ symmetry generated by $\mathcal{A}=(\prod_{l}\sigma^{z}_{l})\mathcal{T}$, where $\mathcal{T}$ is the usual time-reversal transformation.
Despite the fairly low symmetry of the IDWM, numerical calculations suggested that its low-energy spectrum closely resembles that of the spin-$1/2$ Heisenberg chain and is consistent with the SU(2)1 WZW theory [19]. Further numerical evidence supporting the SU(2)1 WZW effective theory is provided in the Supplemental Materials (SM)  [^1] See the Supplemental Materials for details. based on the two-interval mutual information of the ground state [26, 27]. However, the microscopic origin of the emergent SU(2)1 conformal symmetry was not clear.
Figure 1: Schematic illustration of the bond and spin operator indices in $\tilde{H}_{0}$ and $\tilde{H}_{h}$.
In order to bring the IDWM into a more tractable form, we apply the following unitary transformation,
$U_{1}=\prod_{n}\exp\Big{(}\frac{i\pi}{4}(\sigma^{z}_{2n}\sigma^{z}_{2n+1}-\sigma^{z}_{2n})\Big{)},$
which is similar to the self-duality transformation in Eq. (2) but acts on every other bond and site. It maps the IDWM into $U_{1}^{\dagger}H_{0}U_{1}$, which contains only nearest-neighbor interactions 11footnotemark: 1. We then apply another unitary transformation composed of single-site operators,
$\begin{split}U_{2}=&\prod_{l}\sigma^{z}_{l}\prod_{n}R_{3n+1}^{\dagger}\Big{(}[111],\frac{2\pi}{3}\Big{)}R_{3n+2}\Big{(}[111],\frac{2\pi}{3}\Big{)}\\ \cdot&\prod_{k}\sigma^{z}_{6k}\sigma^{x}_{6k+1}\sigma^{x}_{6k+2}\sigma^{y}_{6k+3}\sigma^{y}_{6k+4}\sigma^{z}_{6k+5},\end{split}$
where $R_{l}([111],2\pi/3)=\exp\big{(}i\pi(\sigma^{x}_{l}+\sigma^{y}_{l}+\sigma^{z}_{l})/3\sqrt{3}\big{)}$ is the spin rotation around the $[111]$ axis by $32\pi/$ 3 and permutes the Pauli operators cyclically, $R_{l}^{\dagger}(\sigma^{x}_{l},\sigma^{y}_{l},\sigma^{z}_{l})R_{l}=(\sigma^{y}_{l},\sigma^{z}_{l},\sigma^{x}_{l})$. This maps the IDWM into a deformed Heisenberg chain, $\tilde{H}_{0}=U_{2}^{\dagger}U_{1}^{\dagger}H_{0}U_{1}U_{2}$,
$\tilde{H}_{0}=\sum_{l\in\gamma}(\sigma_{l}^{\alpha}\sigma_{l+1}^{\alpha}+\sigma_{l}^{\beta}\sigma_{l+1}^{\beta}),$ (3)
where $\gamma$ denotes the bond index shown in Fig. 1, which has a three-site periodicity, and $(\alpha,\beta,\gamma)$ form a right-handed frame.
The deformed Heisenberg chain in Eq. (3) has exactly the same form as the Kitaev-Gamma chain in a rotated basis studied in Ref. [28], and has a nonsymmorphic octahedral symmetry, which we explain below. It has a three-site translation symmetry $T_{3}$ ($T_{n}$ denotes the lattice translation by $n$ sites), three global spin flip symmetries $\mathcal{R}(\hat{n},\pi)=\prod_{l}R_{l}(\hat{n},\pi)$ ($\hat{n}=\hat{x}$, $\hat{y}$, $\hat{z}$), and an antiunitary time-reversal symmetry $\mathcal{T}$. Moreover, it is also invariant under the spin-screw translation $\mathcal{R}_{T}T_{1}$ and the spin-flip reflection $\mathcal{R}_{I}I_{0}$, where $\mathcal{R}_{T}=\prod_{l}R_{l}([111],-2\pi/3)$ is a spin rotation, $\mathcal{R}_{I}=\prod_{l}R_{l}([1\bar{1}0],\pi)$ is a spin flip, and $I_{0}$ denotes the site-00-centered lattice reflection. These symmetry operators generate a nonsymmorphic group $G_{0}$, which has a normal subgroup generated by the three-site translation $\langle T_{3}\rangle\cong\mathbb{Z}$, and $G_{0}/\langle T_{3}\rangle\cong O_{h}$, where $O_{h}$ denotes the full octahedral group [28].
It has been shown that the low-energy physics of the deformed Heisenberg chain Eq. (3) is described by the SU(2)1 WZW theory [28]. Let us briefly recapitulate the arguments below. In the non-Abelian bosonization formalism, the spin operators are mapped to the following continuum fields [29, 30],
$\mathbf{S}(x)\simeq\mathbf{J}(x)+\bar{\mathbf{J}}(x)+(-1)^{x}\mathbf{N}(x),$ (4)
where $\mathbf{J}(x)$ and $\bar{\mathbf{J}}(x)$ are the right- and the left-moving current operators, respectively, and $\mathbf{N}(x)$ is the staggered component of the local spin density. The Heisenberg chain is described by the following Hamiltonian density in the continuum limit [29, 30],
$\mathcal{H}_{0}\simeq\frac{2\pi}{3}v(\mathbf{J}\cdot\mathbf{J}+\bar{\mathbf{J}}\cdot\bar{\mathbf{J}})+\lambda\mathbf{J}\cdot\bar{\mathbf{J}},$ (5)
where $v$ is the velocity of low-energy excitations and $\lambda$ is a nonuniversal coupling constant. The first term is exactly the SU(2)1 WZW model, while the second term is an SU(2) invariant perturbation, which is marginally irrelevant for $\lambda<0$. It turns out that these two terms are the only relevant or marginal operators compatible with the nonsymmorphic octahedral symmetry [28]. For example, the current operator bilinear terms that break the SU(2) symmetry are prohibited by $\mathcal{R}_{T}T_{1}$ and the global spin flip symmetries, while the bond dimerization term $\epsilon$ is prohibited by $\mathcal{R}_{T}T_{1}$ and $\mathcal{R}_{I}I_{0}$ symmetries. Therefore, the IDWM has an emergent SU(2)1 conformal symmetry in the low-energy limit and is described by the SU(2)1 WZW theory.
Domain wall transition.— A weak uniform magnetic field in the bulk introduces a nonzero string tension of the $\mathbb{Z}_{2}$ gauge field described by $h\sum_{l}\tau_{l}^{z}$, where $\tau_{l}^{z}$ acts on the bulk spins. While it is an irrelevant perturbation to the bulk topological orders because of the bulk energy gap, it turns out to drive a domain wall transition. Following the derivation of the IDWM, this term is mapped to an Ising interaction in the domain wall model 11footnotemark: 1,
$H_{h}=-h\sum_{l}\sigma^{z}_{l}\sigma^{z}_{l+1}.$
It respects the self-duality symmetry and the antiunitary symmetry, $\mathcal{S}^{\dagger}H_{h}\mathcal{S}=H_{h}$, $\mathcal{A}^{-1}H_{h}\mathcal{A}=H_{h}$.
Applying the same unitary transformations $U_{1}U_{2}$, we find
$\tilde{H}_{h}=h\sum_{l\in\gamma}\sigma_{l}^{\beta}\sigma_{l+1}^{\alpha},$
where we adopt the same notations of bond and spin operator indices illustrated in Fig. 1. While the global spin flip symmetries $\mathcal{R}(\hat{n},\pi)$ ($\hat{n}=\hat{x},\hat{y},\hat{z}$) are explicitly broken, the nonsymmorphic $\mathcal{R}_{T}T_{1}$ and $\mathcal{R}_{I}I_{0}$ and the time-reversal symmetry $\mathcal{T}$ are still preserved. The symmetry group $G_{1}$ generated by these operators satisfies $G_{1}/\langle T_{3}\rangle\cong D_{3h}$.
With the bosonization formula (4) and the operator product expansion (OPE) relations listed in the SM 11footnotemark: 1, $H_{h}$ is mapped to the following Hamiltonian density,
$\mathcal{H}_{h}\simeq\frac{4}{3}h(J_{z}\bar{J}_{y}+\bar{J}_{z}J_{y}+J_{y}\bar{J}_{x}+\bar{J}_{y}J_{x}+J_{x}\bar{J}_{z}+\bar{J}_{x}J_{z}),$
where $\mathbf{J}$ and $\bar{\mathbf{J}}$ terms are omitted because they are conserved currents and can be eliminated by redefining the boson field $\varphi_{s}$ introduced below. This symmetric bilinear term can be simplified by an orthogonal transformation, $\mathbf{M}=V\mathbf{J}$ and $\bar{\mathbf{M}}=V\bar{\mathbf{J}}$, where $V=R(\hat{x},\arctan\sqrt{2})R(\hat{z},\pi/4)$ rotates the $[111]$ axis into the $\hat{z}$ axis, and we find
$\mathcal{H}_{h}\simeq\frac{4}{3}h(2M_{z}\bar{M}_{z}-M_{x}\bar{M}_{x}-M_{y}\bar{M}_{y}).$
It has an emergent U(1) symmetry of rotation around the new $\hat{z}$ axis, which is enhanced from the nonsymmorphic $D_{3h}$ symmetry of the lattice model.
Figure 2: Perturbative RG flow of (a) the domain wall model $\mathcal{H}_{1}=\mathcal{H}_{0}+\mathcal{H}_{h}$ dictated by Eq. (7) and (b) the spin chain $\mathcal{H}_{2}=\mathcal{H}_{1}+\mathcal{H}_{g}$ given by Eq. (8) in the $K_{s}=0$ limit. In (a), the region in orange is the valence-bond solid (VBS) phase of the AF Heisenberg chain [31], which is not realized in the domain wall model.
The perturbed domain wall Hamiltonian $\mathcal{H}_{1}=\mathcal{H}_{0}+\mathcal{H}_{h}$ is given by
$\begin{split}\mathcal{H}_{1}\simeq&\frac{2\pi}{3}v(\mathbf{M}\cdot\mathbf{M}+\bar{\mathbf{M}}\cdot\bar{\mathbf{M}})\\ +&\lambda_{1}(M_{x}\bar{M}_{x}+M_{y}\bar{M}_{y})+\lambda_{2}M_{z}\bar{M}_{z},\end{split}$ (6)
where $\lambda_{1}=\lambda-4h/3$ and $\lambda_{2}=\lambda+8h/3$. The perturbative renormalization group (RG) equations are derived from the OPE relations in the SM 11footnotemark: 1 [32],
$\frac{\mathrm{d}\lambda_{1}}{\mathrm{d}l}=\frac{1}{2\pi}\lambda_{1}\lambda_{2},\quad\frac{\mathrm{d}\lambda_{2}}{\mathrm{d}l}=\frac{1}{2\pi}\lambda_{1}^{2}.$ (7)
The RG flow dictated by Eq. (7) is plotted in Fig. 2 (a). The bare value of $\lambda$ is negative, thus the IDWM with $h=0$ flows along the purple line towards the SU(2)1 fixed point at $\lambda_{1}=\lambda_{2}=0$. The domain wall perturbed by the Ising interaction $H_{h}$ deviates from this line. For $h<0$, the bare parameters satisfy $\lambda_{2}<\lambda_{1}<0$, and flow to the fixed point on the lower half of the vertical axis, which corresponds to a gapless TLL state. For $h>0$, the bare parameters satisfy $\lambda_{1}<\lambda_{2}<0$, and $\lambda_{1}\rightarrow-\infty$ while $\lambda_{2}\rightarrow+\infty$ with the RG flow, which implies a gapped phase.
The nature of the $h>0$ phase can be further clarified with the Abelian bosonization formalism, in which the Hamiltonian is mapped to the sine-Gordon model 11footnotemark: 1,
$\mathcal{H}_{1}\simeq\frac{1}{2}v\big{(}K_{s}\pi_{s}^{2}+K_{s}^{-1}(\partial_{x}\varphi_{s})^{2}\big{)}-\frac{\lambda_{1}}{4\pi^{2}}\cos\sqrt{8\pi}\varphi_{s},$
where $\varphi_{s}$ and $\pi_{s}$ are the boson field and its conjugate momentum, respectively, and $K_{s}\simeq 1-\lambda_{2}/(4\pi v)$ is the bare Luttinger parameter. In the $h>0$ phase, $\lambda_{1}\rightarrow-\infty$ with the RG flow, hence the boson field $\varphi_{s}$ obtains a nonzero expectation value at the ground state $\langle\varphi_{s}\rangle=\pm\sqrt{\pi/8}$. It spontaneously breaks the time-reversal symmetry $\mathcal{T}:\varphi_{s}\mapsto-\varphi_{s}$, and induces an AF order $\langle N_{z}\rangle\propto\sin\sqrt{2\pi}\langle\varphi_{s}\rangle$ of the deformed Heisenberg chain in the $M$-spin current representation. In the original domain wall model representation, this corresponds to the FM order $\langle\sigma^{z}_{l}\rangle\neq 0$, and spontaneously breaks both the antiunitary $\mathbb{Z}_{2}^{T}$ symmetry and the $\mathbb{Z}_{2}$ self-duality symmetry 11footnotemark: 1. This reflects the anomalous nature of the $\mathbb{Z}_{2}$ symmetry: the domain wall must either be gapless or spontaneously break this symmetry.
Figure 3: Numerical results of the domain wall model $H_{1}$. (a) Half-chain EE $S(L/2)$ versus the chain length $L$ in the logarithmic scale in the parameter range $-1.28\leq h\leq 0$. Solid lines are the fitting to $S(L/2)=(c/6)\ln L+a$, which give an estimate of central charges listed in the legend. (b) Squared FM and AF order parameters $m_{\mathrm{FM}}^{2}$ and $m_{\mathrm{AF}}^{2}$ versus $h$ for $L=400$. All calculations are performed with the open boundary condition.
We perform density-matrix renormalization group (DMRG) calculations on the domain wall model $H_{1}=H_{0}+H_{h}$. The half-chain entanglement entropy (EE) $S(L/2)$ in the parameter range $-1.28\leq h\leq 0$ is plotted in Fig. 3 (a). It obeys the formula $S(L/2)=(c/6)\ln L+a$ [33] and gives an estimate of the central charge $c\simeq 1$ as listed in the legend, which is a hallmark of the TLL state. The squared FM and AF order parameters are estimated with the spin correlation functions, $m_{\mathrm{FM}}^{2}=(2/L)\sum_{r=0}^{L/2}\langle\sigma_{L/2-r/2}^{z}\sigma_{L/2+r/2}^{z}\rangle$ and $m_{\mathrm{AF}}^{2}=(2/L)\sum_{r=0}^{L/2}(-1)^{r}\langle\sigma_{L/2-r/2}^{z}\sigma_{L/2+r/2}^{z}\rangle$, and are plotted in Fig. 3 (b). The TLL undergoes a continuous transition at $h\simeq 0$ to the FM order, which is consistent with the RG analysis. However, we find a transition at $h\simeq-1.3$ from the TLL to the AF order, which is not captured by the RG analysis. The abrupt change of the AF order parameter indicates that this is a first-order transition. The AF order spontaneously breaks the time-reversal, the lattice translation and the $\mathbb{Z}_{2}$ self-duality symmetries, hence it is also a manifestation of the anomaly of $\mathbb{Z}_{2}$ symmetry.
Gapless domain wall as 1D topological transition.— In a holographic construction [22, 23], topological QCPs between SPT phases can be realized as gapless domain wall states between topological phases in one-higher dimensions. In this section, we demonstrate this holographic scenario with the gapless domain wall of $\mathbb{Z}_{2}$ topological orders. We generalize the domain wall model $H_{1}$ and introduce an extra tuning parameter $g$,
$H_{2}=-\sum_{l}\big{(}(1+g)\sigma^{z}_{l-1}\sigma^{x}_{l}\sigma^{z}_{l+1}+(1-g)\sigma^{y}_{l}+h\sigma^{z}_{l}\sigma^{z}_{l+1}\big{)}.$
The $g$-term changes sign under the self-duality transformation, hence it explicitly breaks the anomalous $\mathbb{Z}_{2}$ symmetry and cannot be realized on the domain wall of $\mathbb{Z}_{2}$ topological orders. Nonetheless, the antiunitary $\mathbb{Z}_{2}^{T}$ symmetry $\mathcal{A}=\prod_{l}\sigma_{l}^{z}\mathcal{T}$ is still preserved.
[…]
Applying the above unitary transformations $U_{1}U_{2}$, the $g$-term is mapped to a bond-alternating interaction,
$\tilde{H}_{g}=-g\sum_{l\in\gamma}(-1)^{l}(\sigma_{l}^{\alpha}\sigma_{l+1}^{\alpha}+\sigma_{l}^{\beta}\sigma_{l+1}^{\beta}).$
This term explicitly breaks the nonsymmorphic $\mathcal{R}_{T}T_{1}$ and $\mathcal{R_{I}}I_{0}$ symmetries, hence allows the bond dimerization operator $\epsilon\propto\cos\sqrt{2\pi}\varphi_{s}$ in the effective theory,
$\mathcal{H}_{g}\simeq\frac{g\gamma}{\pi}\cos\sqrt{2\pi}\varphi_{s},$
where $\gamma\neq 0$ is a nonuniversal constant 11footnotemark: 1. The scaling dimension of the $\mathcal{H}_{g}$ term is $1/2$ in the SU(2)1 WZW theory, thereby it is a relevant perturbation and drives the IDWM immediately into the $\mathbb{Z}_{2}^{T}$-SPT phase for $g>0$ and into the trivial phase for $g<0$. Therefore, the IDWM is a 1D topological QCP between the $\mathbb{Z}_{2}^{T}$-SPT and the trivial phases.
Figure 4: (a) Quantum phase diagram of $H_{2}$. The purple segment on the vertical axis indicates the TLL phase of the domain wall model. Phase boundaries are extracted with the crossing of $\Delta L^{z}$ evaluated with different lattice sizes, for which an example at $h=0.3$ is shown in (b).
The spin chain $H_{2}$ is described by the two-frequency sine-Gordon model $\mathcal{H}_{2}=\mathcal{H}_{1}+\mathcal{H}_{g}$. The perturbative RG equations are given by 11footnotemark: 1
$\frac{\mathrm{d}K_{s}}{\mathrm{d}l}=-K_{s}^{2}(4\pi^{2}\tilde{\lambda}_{1}^{2}+\pi^{2}\tilde{g}^{2}),$ (8a)
$\frac{\mathrm{d}\tilde{\lambda}_{1}}{\mathrm{d}l}=(2-2K_{s})\tilde{\lambda}_{1}+\frac{1}{2}\pi\tilde{g}^{2},$ (8b)
$\frac{\mathrm{d}\tilde{g}}{\mathrm{d}l}=\Big{(}2-\frac{1}{2}K_{s}\Big{)}\tilde{g}+\pi\tilde{\lambda}_{1}\tilde{g},$ (8c)
where the rescaled parameters $\tilde{\lambda}_{1}=\lambda_{1}/(8\pi^{2}v)$ and $\tilde{g}=g\gamma/(2\pi^{2}v)$ are introduced for convenience. For $g=0$, the RG equations are consistent with Eq. (7) in the perturbative regime with $\lambda_{1,2}\ll 1$. Starting from the bare parameters $\tilde{\lambda}_{1}<0$ and $K_{s}>0$, the RG flow ends up in either the TLL phase ($K_{s}\rightarrow K_{s}^{*}\geq 1$, $\tilde{\lambda}_{1}\rightarrow 0$) or the FM order ($K_{s}\rightarrow 0$, $\tilde{\lambda}_{1}\rightarrow-\infty$). The $g$-term significantly changes the RG flow. In particular, the TLL state with $K_{s}^{*}<4$ is destabilized by the relevant $g$-term, which drives $K_{s}$ to zero. Focusing on this regime and setting $K_{s}=0$ in the RG equations, we plot the RG flow of $\tilde{\lambda}_{1}$ and $\tilde{g}$ in Fig. 2 (b), where three gapped phases are illustrated: the FM order ($\tilde{g}\rightarrow 0$, $\tilde{\lambda}_{1}\rightarrow-\infty$) spontaneously breaks the $\mathbb{Z}_{2}^{T}$ symmetry, while the other two phases ($\tilde{\lambda}_{1}\rightarrow+\infty$, $\tilde{g}\rightarrow\pm\infty$) are disordered and correspond to the $\mathbb{Z}_{2}^{T}$-SPT phase and trivial phase, respectively. There is a continuous transition between the FM order and either of the disordered phases, which, according to the nonperturbative analysis of the two-frequency sine-Gordon model, belongs to the 2D Ising universality class [37, 38], and thus is smoothly connected with the exactly solvable QCPs of the quantum Ising chains discussed above. In addition, according to Eq. (8), the TLL phase with $K_{s}^{*}>4$ should be stable against the $g$-term perturbation. However, this is not realized in the lattice model with a nonzero $g$-term according to the following numerical results. This also reflects the inadequacy of the perturbative RG analysis.
The quantum phase diagram is obtained with DMRG and plotted in Fig. 4 (a). It is symmetric about the vertical axis, because the $g$-term flips sign under the unitary self-duality transformation. The phase boundaries are extracted with the crossing of $\Delta L^{z}$, where $\Delta$ is the energy gap, and $z=1$ is the dynamical exponent. $\Delta L^{z}$ is scale invariant at the QCP, thus its values for different lattice sizes should cross at the QCP, which is exemplified in Fig. 4 (b). This model has four phases: the FM and the AF ordered phases, and the $\mathbb{Z}_{2}^{T}$-SPT and the trivial gapped phases. The ordered phases are separated from the disordered phases by Ising critical lines. Moreover, a direct topological transition between the $\mathbb{Z}_{2}^{T}$-SPT and the trivial phases takes place at the purple segment in the vertical axis, which corresponds to the gapless TLL state on the domain wall of $\mathbb{Z}_{2}$ topological orders. Therefore, this model realizes the holographic scenario of topological transitions.
Conclusion.— We have theoretically investigated the IDWM of $\mathbb{Z}_{2}$ topological orders and its generalization. We found the equivalence of the IDWM and the deformed Heisenberg chain with a nonsymmorphic octahedral symmetry, and thus unveiled the microscopic origin of the emergent SU(2)1 conformal symmetry. The bulk magnetic field induces domain wall transitions from the TLL phase to the FM and the AF orders. The domain wall is either gapless or symmetry-breaking, which reflects the $\mathbb{Z}_{2}$ symmetry anomaly. Moreover, the TLL state is a holographic realization of 1D topological QCP between the $\mathbb{Z}_{2}^{T}$-SPT and trivial phases as the gapless domain wall of 2D topological orders. Therefore, our work provides a valuable model demonstrating the rich physics on the domain wall of topological orders.
It is desirable to generalize this work and further explore the anomalous symmetry and its consequences on the boundary and domain walls of topological phases with concrete lattice models. The interplay of the bulk topological orders with the boundary and domain wall states also deserves further study, including the generalized symmetry generated by bulk topological excitations and defects acting on the boundary and domain wall states [39], and novel boundary and domain wall critical behavior induced by the interactions of the gapless boundary and domain wall states with the bulk states at the bulk QCP.
Acknowledgements.
We are grateful to helpful discussions with Zheng-Cheng Gu and Chao Xu. Part of the numerical simulations were carried out with the ITensor package [40, 41]. This work was supported by the National Natural Science Foundation of China (12174387 and 12304182), the Chinese Academy of Sciences (YSBR-057 and JZHKYPT-2021-08), and the Innovative Program for Quantum Science and Technology (2021ZD0302600).